Evaluate the\\(\\frac{d^{3}f3{dx^{3}}\\ of \\(f(x)=sin(x)\\)
1
Expert's answer
2019-09-09T11:43:25-0400

Solution. Find the first derivative of the function f(x)

"\frac{df} {dx}=(sin(x))'=cos(x)"

Find the second derivative of the function f(x)


"\frac{d^2f} {dx^2}=(cos(x))'=-sin(x)"

Find the third derivative of the function f(x)


"\frac{d^3f} {dx^3}=(-sin(x))'=-cos(x)"

Answer. - cos(x).